OpenStudy (anonymous):

Simplify the rational expression. State any excluded values. x-2/x^2+3x-10

OpenStudy (anonymous):

you have to factor the denominator x^2 + 3x - 10 do you know how to factor it ?

OpenStudy (anonymous):

No can you show me how

OpenStudy (anonymous):

ax^2 + bx + c you have to split the 'b' into two terms say b1 and b2 ax^2 + b1x + b2x +c your b1 and b2 have to satisfy the conditions : b1 + b2 =2 and b1 * b2 = a * c in this case we can choose b1 to be 5 and b2 to be -2 so : 5 - 2 = 3 and 5*-2 = -10

OpenStudy (anonymous):

then we will have x^2 + 5x - 2x - 10 in the denominator now can you factor it ?

OpenStudy (anonymous):

im trying to understand it

OpenStudy (anonymous):

splitting 'b' is a method that will help you to eventually factor it .. it is important to understand this method

OpenStudy (anonymous):

How I describe factorization is we first look at the constant only. And we write down or think about all the numbers, when multiplied will give us this value.

OpenStudy (anonymous):

So for -10 it's: -1,10 1,-10 2,-5 -2,5

OpenStudy (anonymous):

Now we look at the coefficient of the middle term and think what pair of numbers when i add them will give me this coefficient

OpenStudy (anonymous):

The coefficient is 3 so if we add the pair -2,5 we get 3

OpenStudy (anonymous):

Now we know it must be in the form (x-2)(x+5) Does this help?

OpenStudy (anonymous):

yes im still trying to work it though

OpenStudy (anonymous):

After you factor the denominator it should look like this \[\frac{ x-2 }{ (x-2)(x+5) }\]Do you see anything that can be simplified from here?

OpenStudy (anonymous):

My description of factoring is the same as @Coolsector 's just with more words

OpenStudy (anonymous):

umm no?

OpenStudy (anonymous):

cant you see a term that it is in the denominator and in the nominator ?

OpenStudy (anonymous):

if you substitute x-2=y \[\frac{ y }{ y(x+5) }\]Now what do you see?

OpenStudy (anonymous):

x-2/x^2+3x-10\[\frac{ x - 2 }{ x^2 + 3x - 10 } = \frac{ x - 2 }{ (x - 2)(x + 5) } = \frac{ 1 }{ x + 5 }\]

OpenStudy (anonymous):

How about "8"=x-2 \[\frac{ 8 }{ 8(x+5) }\]

OpenStudy (anonymous):

plz don't give the answer @KKJ it helps them more to work through it

OpenStudy (anonymous):

The major problem here is how to factorize the quadratic function

OpenStudy (anonymous):

thats how i was working it out kkj but i was confused Thank you

OpenStudy (anonymous):

Thanks guys!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[x^2- 36\] is difference of two squares

OpenStudy (anonymous):

@kkj its on the new on i just posted not this one lol

OpenStudy (anonymous):

\[\frac{ x^2-36 }{ 6 - x } = \frac{ (x - 6)(x + 6) }{ -1(x - 6) } = \frac{ x + 6 }{ -1 } = -x - 6\]

OpenStudy (anonymous):

ty